# Constructing Concise Characteristic Samples for Acceptors of Omega Regular Languages

### v3

@@ 13. An inclusion algorithm for DRAs
Algorithm 1 $\mathbf{SubInc}^{DRA}$
Two DRAs $\mathcal{R}_{1}=(\mathcal{M}_{1},\alpha_{1})$ and $\mathcal{R}_{2}=(\mathcal{M}_{2},\alpha_{2})$ in singleton normal form, where $\alpha_{1}=\{(q^{\prime},B^{\prime})\}$ and $\alpha_{2}=\{(q_{1}^{\prime\prime},B_{1}^{\prime\prime}),\ldots,(q_{k}^{\prime\prime},B_{k}^{\prime\prime})\}$, and a set $S$ of states of $\mathcal{M}=\mathcal{M}_{1}\times\mathcal{M}_{2}$
@@
           Call $\mathbf{SubInc}^{DRA}$ recursively with $\mathcal{R}_{1}$, $\mathcal{R}_{2}$, and $S^{\prime\prime}$
@@
The algorithm $\mathbf{SubInc}^{DRA}$ takes as input two DRAs $\mathcal{R}_{1}=(\mathcal{M}_{1},\alpha_{1})$ and $\mathcal{R}_{2}=(\mathcal{M}_{2},\alpha_{2})$ in singleton normal form, where $\alpha_{1}$ consists of a single pair $(q^{\prime},B^{\prime})$. It also takes as input a subset $S$ of the state set of the product automaton $\mathcal{M}=\mathcal{M}_{1}\times\mathcal{M}_{2}$. The problem it solves is to determine whether there exists an $\omega$-word $u(v)^{\omega}$ with ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega})\subseteq S$ such that $u(v)^{\omega}\in{\llbracket}\mathcal{R}_{1}{\rrbracket}\setminus{\llbracket}\mathcal{R}_{2}{\rrbracket}$. If there is such a word, the algorithm returns one with the $\preceq$-largest value of ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega})$, and otherwise, it returns “none”.
@@
For $i=1,2$ let $\mathcal{R}_{i}=(\mathcal{M}_{i},\alpha_{i})$ be a DRA in singleton normal form. Assume $\alpha_{1}=\{(q^{\prime},B^{\prime})\}$ and $\alpha_{2}=\{(q_{1}^{\prime\prime},B_{1}^{\prime\prime}),\ldots,(q_{k}^{\prime\prime},B_{k}^{\prime\prime})\}$. Let $\mathcal{M}=\mathcal{M}_{1}\times\mathcal{M}_{2}$, and let $S$ be a subset of the states of $\mathcal{M}$. Then with inputs $\mathcal{R}_{1}$, $\mathcal{R}_{2}$, and $S$, the algorithm $\mathbf{SubInc}^{DRA}$ runs in polynomial time and returns $u(v)^{\omega}\in{\llbracket}\mathcal{R}_{1}{\rrbracket}\setminus{\llbracket}\mathcal{R}_{2}{\rrbracket}$ with the $\preceq$-largest value of ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega})$ contained in $S$, if such exists, else it returns “none”.
@@
We now show by induction on the recursive calls that if $w$ is any $\omega$-word such that ${{\textsl{inf}}}_{\mathcal{M}}(w)\subseteq S$, $(\mathcal{M}_{1},\{(q^{\prime},B^{\prime})\})$ accepts $w$, and $\mathcal{R}_{2}$ rejects $w$, then $\mathbf{SubInc}^{DRA}$ returns a witness $u(v)^{\omega}$ such that ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega})$ is at least as large as ${{\textsl{inf}}}_{\mathcal{M}}(w)$ in the $\preceq$-ordering. Let $D={{\textsl{inf}}}_{\mathcal{M}}(w)$. Then $D$ is a reachable SCC of $\mathcal{M}$ such that $q^{\prime}\in\pi_{1}(D)$, $B^{\prime}\cap\pi_{1}(D)=\emptyset$, and for no $j\in[1..k]$ do we have $q_{j}^{\prime\prime}\in\pi_{2}(D)$ and $B_{j}^{\prime\prime}\cap\pi_{2}(D)=\emptyset$. Then $D\subseteq S^{\prime}$ because $B^{\prime}\cap\pi_{1}(D)=\emptyset$. Thus, $D$ must be a subset of exactly one of the elements $C$ of $\textit{maxSCCs}(S^{\prime})$. Then $C$ is reachable, $q^{\prime}\in\pi_{1}(C)$, and $B^{\prime}\cap\pi_{1}(C)=\emptyset$ (because $C$ is a subset of $S^{\prime}$).
@@
Otherwise, the set $J=\{q_{j}^{\prime\prime}\mid j\in[1..k],B_{j}^{\prime\prime}\cap\pi_{2}(C)=\emptyset\}$ is non-empty, and the algorithm removes from $C$ all the states $(q_{1},q_{2})$ such that $q_{2}\in J$ to form the set $S^{\prime\prime}$. Because $D\subseteq C$, if $B_{j}^{\prime\prime}\cap\pi_{2}(C)=\emptyset$, then also $B_{j}^{\prime\prime}\cap\pi_{2}(D)=\emptyset$. Thus, if for any $q_{j}^{\prime\prime}\in J$ we have $q_{j}^{\prime\prime}\in\pi_{2}(D)$, this would violate the assumption that $\mathcal{R}_{2}$ rejects $w$. Hence, $D\subseteq S^{\prime\prime}$, and by the inductive assumption on the recursive calls, the recursive call to $\mathbf{SubInc}^{DRA}$ returns a witness $u(v)^{\omega}$ such that ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega})$ is at least as large in the $\preceq$-ordering as ${{\textsl{inf}}}_{\mathcal{M}}(w)$. Because the top-level algorithm returns $u(v)^{\omega}$ to maximize ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega})$ with respect to $\preceq$, it will be at least as large as ${{\textsl{inf}}}_{\mathcal{M}}(w)$.
@@
There are polynomial time algorithms to solve the inclusion and equivalence problems for two DRAs (resp. DSAs) $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$. In the case of non-inclusion or non-equivalence, these algorithms return a witness $u(v)^{\omega}$ with the $\preceq$-largest value of ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega})$, where $\mathcal{M}$ is the product of the automata of $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$.
@@
It suffices to consider just DRAs, by Claim 5. Given two DRAs $\mathcal{R}_{1}=(\mathcal{M}_{1},\alpha_{1})$ and $\mathcal{R}_{2}=(\mathcal{M}_{2},\alpha_{2})$, we may assume they are in singleton normal form. Then for each pair $(q_{i},B_{i})$ in $\alpha_{1}$, we call $\mathbf{SubInc}^{DRA}$ with inputs $(\mathcal{M}_{1},\{(q_{i},B_{i})\})$, $\mathcal{R}_{2}$, with $S$ equal to the whole state set of $\mathcal{M}$. If all of these calls return “none”, then ${\llbracket}\mathcal{R}_{1}{\rrbracket}$ is a subset of ${\llbracket}\mathcal{R}_{2}{\rrbracket}$, and the answer returned is “yes”. Otherwise, one or more calls return a witness, and $u(v)^{\omega}$ is returned such that ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega})$ is $\preceq$-largest among the witnesses returned by the calls. The running time and correctness follow from the running time and correctness guarantees of $\mathbf{SubInc}^{DRA}$. ∎
@@ 14.1. Reduction of DMA inclusion to DBA/DMA inclusion
Intuitively, for an input from $L(\mathcal{M},F,q)$, in $B(\mathcal{M},F,q)$ the states $q_{i}$ are visited in a repeating cyclic order: $q_{0},q_{1},\ldots,q_{m-1}$, and the meaning of the state $r_{i,j}$ is that at this point in the input, $\mathcal{M}$ would be in state $q_{i}$, and the machine $B(\mathcal{M},F,q)$ is waiting for a transition that would arrive at state $q_{j}$ in $\mathcal{M}$, in order to proceed to state $q_{j}$ in $B(\mathcal{M},F,q)$. [^9] This construction is reminiscent of the construction transforming a generalized Büchi into a Büchi automaton [Var08, Cho74], by considering each state in $F$ as a singleton set of a generalized Büchi, but here we need to send transitions to states outside $F$ to a sink state. An example of the construction is shown in Fig. 4; the dead state and unreachable states are omitted for clarity.
@@ 14.2. A DBA/DMA inclusion algorithm
Then $\pi_{2}(C^{\prime})\subseteq F^{\prime}$ but because $\pi_{2}(C^{\prime})\not\in\mathcal{F}$, it must be that $\pi_{2}(C^{\prime})\neq F^{\prime}$. Thus, there exists some $q\in F^{\prime}$ that is not in $\pi_{2}(C^{\prime})$. When the algorithm considers this $F^{\prime}$ and $q$, then because $\pi_{2}(C^{\prime})\subseteq F^{\prime}\setminus\{q\}$, $C^{\prime}$ is contained $R$ and therefore is a subset of a unique SCC $D$ in $\textit{maxSCCs}(R)$.
@@
Combining Thm. 14.4, Thm. 14.5, and the reduction of equivalence to inclusion, we have the following.
@@ 15. Computing the automaton $\mathcal{M}_{\sim_{L}}$
An acceptor $\mathcal{A}=\langle\Sigma,Q,q_{\iota},\delta,\alpha\rangle$ of type DBA, DCA, DPA, DRA, DSA, or DMA
A deterministic automaton $\mathcal{M}$ isomorphic to $\mathcal{M}_{\sim_{L}}$, where $L={\llbracket}\mathcal{A}{\rrbracket}$
@@
To test whether a given DBA (resp., DCA, DPA, DRA, DSA, DMA) $\mathcal{A}$ is an IBA (resp., ICA, IPA, IRA, ISA, IMA), we run the $\mathbf{RightCon}$ algorithm on $\mathcal{A}$ and test the returned automaton $\mathcal{M}$ for isomorphism with the automaton of $\mathcal{A}$. If they are isomorphic, then $\mathcal{A}$ is an IBA (resp. ICA, IPA, IRA, ISA, IMA), otherwise it is not. This proves the following.
@@ 16. Testing membership in $\mathbb{IXA}$
In the previous section we showed that there is a polynomial time algorithm to test whether a given DBA $\mathcal{B}$ is an IBA. However, we can also ask the following harder question. Given a DBA $\mathcal{B}$ that is not an IBA, is ${\llbracket}\mathcal{B}{\rrbracket}\in\mathbb{IBA}$, that is, does there exist an IBA $\mathcal{B}^{\prime}$ such that ${\llbracket}\mathcal{B}^{\prime}{\rrbracket}={\llbracket}\mathcal{B}{\rrbracket}$? This section shows that there are such polynomial time algorithms for DBAs, DCAs, DPAs, DRAs, DSAs, and DMAs. The algorithms first compute the right congruence automaton $\mathcal{M}=\mathcal{M}_{\sim_{L}}$, where $L={\llbracket}\mathcal{A}{\rrbracket}$, and then attempts to construct an acceptance condition $\alpha$ of the appropriate type such that ${\llbracket}(\mathcal{M},\alpha){\rrbracket}=L$.
@@ 16.1. Testing membership in $\mathbb{IBA}$
We describe the algorithm $\mathbf{TestInIB}$ that takes as input a DBA $\mathcal{B}$ and returns an IBA accepting ${\llbracket}\mathcal{B}{\rrbracket}$ if ${\llbracket}\mathcal{B}{\rrbracket}\in\mathbb{IBA}$, and otherwise returns “no”. By Claim 4, the case of a DCA is reduced to that of a DBA.
@@
A DBA $\mathcal{B}$
If ${\llbracket}\mathcal{B}{\rrbracket}\in\mathbb{IBA}$ then return an IBA accepting ${\llbracket}\mathcal{B}{\rrbracket}$, else return “no”
@@ A DBA $\mathcal{B}$
The algorithm calls the $\mathbf{RightCon}$ algorithm, and also the inclusion and equivalence algorithms from Thm. 12.3, which run in polynomial time in the size of $\mathcal{B}$. If the algorithm returns an acceptor, it is an IBA accepting ${\llbracket}\mathcal{B}{\rrbracket}$.
@@ 16.2. Testing membership in $\mathbb{IPA}$
A DPA $\mathcal{P}$
If ${\llbracket}\mathcal{P}{\rrbracket}\in\mathbb{IPA}$ then return an IPA accepting ${\llbracket}\mathcal{P}{\rrbracket}$, else return “no”
@@ A DPA $\mathcal{P}$
The algorithm calls the $\mathbf{RightCon}$ algorithm and the inclusion and equivalence algorithms for DPAs from Thm. 12.2, which run in polynomial time in the size of $\mathcal{P}$. Below we show that each while loop terminates after at most $|Q|$ iterations. If the algorithm returns an acceptor, then the acceptor is an IPA accepting ${\llbracket}\mathcal{P}{\rrbracket}$.
@@ 16.3. Testing membership in $\mathbb{IRA}$
We describe the algorithm $\mathbf{TestInIR}$ that takes as input a DRA $\mathcal{R}$ and returns an IRA accepting ${\llbracket}\mathcal{R}{\rrbracket}$ if ${\llbracket}\mathcal{R}{\rrbracket}\in\mathbb{IRA}$, and otherwise returns “no”. By Claim 5, the case of a DSA is reduced to that of a DRA.
@@
Given a pair $(q,B)\in\alpha_{1}$ the construction of the initially empty acceptance condition $\alpha^{\prime}$ proceeds as follows. Let $C_{0}$ be the maximum SCC of $\mathcal{M}_{1}$ that contains $q$ and excludes $B$. If $C_{0}$ is empty, then $(\mathcal{M}_{1},\{(q,B)\})$ does not accept any words, and the empty condition $\alpha^{\prime}$ suffices. If $C_{0}$ is nonempty, then there is an element $u(v)^{\omega}$ of ${\llbracket}(\mathcal{M}_{1},\{(q,B)\}{\rrbracket}$ such that $C_{0}={{\textsl{inf}}}_{\mathcal{M}_{1}}(u(v)^{\omega})$ and there is a pair $(q_{0},B_{0})$ in $\alpha_{2}$ such that $q_{0}\in f(C_{0})$ and $B_{0}\cap f(C_{0})=\emptyset$. We add the pair $(q_{0},Q\setminus f(C_{0}))$ to $\alpha^{\prime}$ and note that by the argument in the preceding paragraph, ${\llbracket}(\mathcal{M},\alpha^{\prime}){\rrbracket}\subseteq{\llbracket}\mathcal{R}{\rrbracket}$.
If ${\llbracket}(\mathcal{M}_{1},\{(q,B)\}{\rrbracket}\subseteq{\llbracket}(\mathcal{M},\alpha^{\prime}){\rrbracket}$ then $\alpha^{\prime}$ is the desired acceptance condition. If not, there exists a word $u(v)^{\omega}$ such that for $C={{\textsl{inf}}}_{\mathcal{M}_{1}}(u(v)^{\omega})$ we have $q\in C$ and $C\cap B=\emptyset$, but either $q_{0}\not\in f(C)$ or $f(C)\cap(Q\setminus f(C_{0}))\neq\emptyset$. Because $C_{0}$ is the maximum SCC of $\mathcal{M}_{1}$ containing $q$ and excluding $B$, we have $C\subseteq C_{0}$, so $f(C)\subseteq f(C_{0})$ and therefore $q_{0}\not\in f(C)$. Let $C_{1}$ be the maximum SCC $C$ of $\mathcal{M}_{1}$ such that $C\subseteq C_{0}$, $q\in C$, and $q_{0}\not\in f(C)$. This is not empty, so there is a word $u^{\prime}(v^{\prime})^{\omega}$ such that $C_{1}={{\textsl{inf}}}_{\mathcal{M}_{1}}(u^{\prime}(v^{\prime})^{\omega})$, which is in ${\llbracket}\mathcal{R}{\rrbracket}$ because it satisfies $(q,B)$. Thus there exists a pair $(q_{1},B_{1})$ in $\alpha_{2}$ that is satisfied by $u^{\prime}(v^{\prime})^{\omega}$, and we add the pair $(q_{1},Q\setminus f(C_{1}))$ to the acceptance condition $\alpha^{\prime}$. As above, we have ${\llbracket}(\mathcal{M},\alpha^{\prime}){\rrbracket}\subseteq{\llbracket}\mathcal{R}{\rrbracket}$.
@@
A DRA $\mathcal{R}=(\mathcal{M}_{1},\alpha_{1})$ in singleton normal form with $|\alpha_{1}|=m$
If ${\llbracket}\mathcal{R}{\rrbracket}\in\mathbb{IRA}$ then return an IRA accepting ${\llbracket}\mathcal{R}{\rrbracket}$, else return “no”
@@
Assume the input is DRA $\mathcal{R}=(\mathcal{M}_{1},\alpha_{1})$ in singleton normal form with $|\alpha_{1}|=m$. The algorithm $\mathbf{TestInIR}$ computes the right congruence automaton $\mathcal{M}$ of ${\llbracket}\mathcal{R}{\rrbracket}$, which has $n$ states, at most the number of states of $\mathcal{R}$. The main loop of the algorithm is executed at most $mn^{3}$ times, and each execution makes calls to the inclusion and equivalence algorithms for DRAs, and runs in time polynomial in the size of $\mathcal{R}$, so the overall running time of $\mathbf{TestInIR}$ is polynomial in the size of $\mathcal{R}$.

### v4

@@ 13. An inclusion algorithm for DRAs
[…]
Two DRAs $\mathcal{R}_{1}=(\mathcal{M}_{1},\alpha_{1})$ and $\mathcal{R}_{2}=(\mathcal{M}_{2},\alpha_{2})$ in singleton normal form, where $\alpha_{1}=\{(q^{\prime},B^{\prime})\}$ and $\alpha_{2}=\{(q_{1}^{\prime\prime},B_{1}^{\prime\prime}),\ldots,(q_{k}^{\prime\prime},B_{k}^{\prime\prime})\}$, and a set $S$ of states of $\mathcal{M}=\mathcal{M}_{1}\times\mathcal{M}_{2}$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
There are polynomial time algorithms to solve the inclusion and equivalence problems for two DRAs (resp. DSAs) $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$. In the case of non-inclusion or non-equivalence, these algorithms return a witness $u(v)^{\omega}$ with the $\preceq$-largest value of ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega})$, where $\mathcal{M}$ is the product of the automata $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$.
@@
[…]
@@ 14.1. Reduction of DMA inclusion to DBA/DMA inclusion
Intuitively, for an input from $L(\mathcal{M},F,q)$, in $B(\mathcal{M},F,q)$ the states $q_{i}$ are visited in a repeating cyclic order: $q_{0},q_{1},\ldots,q_{m-1}$, and the meaning of the state $r_{i,j}$ is that at this point in the input, $\mathcal{M}$ would be in state $q_{i}$, and the machine $B(\mathcal{M},F,q)$ is waiting for a transition that would arrive at state $q_{j}$ in $\mathcal{M}$, in order to proceed to state $q_{j}$ in $B(\mathcal{M},F,q)$. [^9] This construction is reminiscent of the construction transforming a generalized Büchi into a Büchi automaton [Var08, Cho74], by considering each state in $F$ as a singleton set of a generalized Büchi, but here we need to send transitions to states outside $F$ to a sink state. An example of the construction is shown in Figure 4; the dead state and unreachable states are omitted for clarity.
@@ 14.2. A DBA/DMA inclusion algorithm
Then $\pi_{2}(C^{\prime})\subseteq F^{\prime}$ but because $\pi_{2}(C^{\prime})\not\in\mathcal{F}$, it must be that $\pi_{2}(C^{\prime})\neq F^{\prime}$. Thus, there exists some $q\in F^{\prime}$ that is not in $\pi_{2}(C^{\prime})$. When the algorithm considers this $F^{\prime}$ and $q$, then because $\pi_{2}(C^{\prime})\subseteq F^{\prime}\setminus\{q\}$, $C^{\prime}$ is contained in $R$ and therefore is a subset of a unique SCC $D$ in $\textit{maxSCCs}(R)$.
@@
Combining Theorem 14.4, Theorem 14.5, and the reduction of equivalence to inclusion, we have the following.
@@ 15. Computing the automaton $\mathcal{M}_{\sim_{L}}$
An acceptor $\mathcal{A}=\langle\Sigma,Q,q_{\iota},\delta,\alpha\rangle$ of type DBA, DCA, DPA, DRA, DSA, or DMA.
A deterministic automaton $\mathcal{M}$ isomorphic to $\mathcal{M}_{\sim_{L}}$, where $L={\llbracket}\mathcal{A}{\rrbracket}$.
@@
To test whether a given DBA (resp., DCA, DPA, DRA, DSA, DMA) $\mathcal{A}$ is an IBA (resp., ICA, IPA, IRA, ISA, IMA), we run the $\mathbf{RightCon}$ algorithm on $\mathcal{A}$ and test the returned automaton $\mathcal{M}$ for isomorphism with the automaton of $\mathcal{A}$. (Note that isomorphism can be checked in polynomial time by gradually constructing a map $h$ between states of $\mathcal{A}$ to states of $\mathcal{M}$. Initially $h(q_{\iota})=q^{\prime}_{\iota}$. Assume $h(q)=q^{\prime}$ and $\delta(q,\sigma)=p$, $\delta(q^{\prime},\sigma)=p^{\prime}$. If $h(p)$ is defined and is different from $p^{\prime}$ return “non-isomorphic”, otherwise set $h(p)=p^{\prime}$. This is repeated until all states are mapped and have been tested with respect to each letter of the alphabet.) If they are isomorphic, then $\mathcal{A}$ is an IBA (resp. ICA, IPA, IRA, ISA, IMA), otherwise it is not. This proves the following.
@@ 16. Testing membership in $\mathbb{IXA}$
In the previous section we showed that there is a polynomial time algorithm to test whether a given DBA $\mathcal{B}$ is an IBA. However, we can also ask the following harder question. Given a DBA $\mathcal{B}$ that is not an IBA, is ${\llbracket}\mathcal{B}{\rrbracket}\in\mathbb{IBA}$, that is, does there exist an IBA $\mathcal{B}^{\prime}$ such that ${\llbracket}\mathcal{B}^{\prime}{\rrbracket}={\llbracket}\mathcal{B}{\rrbracket}$? This section shows that there are such polynomial time algorithms for DBAs, DCAs, DPAs, DRAs, DSAs and DMAs. The algorithms first compute the right congruence automaton $\mathcal{M}=\mathcal{M}_{\sim_{L}}$, where $L={\llbracket}\mathcal{A}{\rrbracket}$, and then attempt to construct an acceptance condition $\alpha$ of the appropriate type such that ${\llbracket}(\mathcal{M},\alpha){\rrbracket}=L$.
@@ 16.1. Testing membership in $\mathbb{IBA}$
We describe the algorithm $\mathbf{TestInIB}$ that takes as input a DBA $\mathcal{B}$ and returns an IBA accepting ${\llbracket}\mathcal{B}{\rrbracket}$ if ${\llbracket}\mathcal{B}{\rrbracket}\in\mathbb{IBA}$, and otherwise returns “no”. By Claim 2.2 (4), the case of a DCA is reduced to that of a DBA.
@@
A DBA $\mathcal{B}$.
If ${\llbracket}\mathcal{B}{\rrbracket}\in\mathbb{IBA}$ then return an IBA accepting ${\llbracket}\mathcal{B}{\rrbracket}$, else return “no”.
@@
The algorithm calls the $\mathbf{RightCon}$ algorithm, and also the inclusion and equivalence algorithms from Theorem 12.3, which run in polynomial time in the size of $\mathcal{B}$. If the algorithm returns an acceptor, it is an IBA accepting ${\llbracket}\mathcal{B}{\rrbracket}$.
@@ 16.2. Testing membership in $\mathbb{IPA}$
A DPA $\mathcal{P}$.
If ${\llbracket}\mathcal{P}{\rrbracket}\in\mathbb{IPA}$ then return an IPA accepting ${\llbracket}\mathcal{P}{\rrbracket}$, else return “no”.
@@
The algorithm calls the $\mathbf{RightCon}$ algorithm and the inclusion and equivalence algorithms for DPAs from Theorem 12.2, which run in polynomial time in the size of $\mathcal{P}$. Below we show that each while loop terminates after at most $|Q|$ iterations. If the algorithm returns an acceptor, then the acceptor is an IPA accepting ${\llbracket}\mathcal{P}{\rrbracket}$.
@@ 16.3. Testing membership in $\mathbb{IRA}$
We describe the algorithm $\mathbf{TestInIR}$ that takes as input a DRA $\mathcal{R}$ and returns an IRA accepting ${\llbracket}\mathcal{R}{\rrbracket}$ if ${\llbracket}\mathcal{R}{\rrbracket}\in\mathbb{IRA}$, and otherwise returns “no”. By Claim 2.2 (5), the case of a DSA is reduced to that of a DRA.
@@
Given a pair $(q,B)\in\alpha_{1}$ the construction of the initially empty acceptance condition $\alpha^{\prime}$ proceeds as follows. Let $C_{0}$ be the maximum SCC of $\mathcal{M}_{1}$ that contains $q$ and contains no element of $B$. If $C_{0}$ is empty, then $(\mathcal{M}_{1},\{(q,B)\})$ does not accept any words, and the empty condition $\alpha^{\prime}$ suffices. If $C_{0}$ is nonempty, then there is an element $u(v)^{\omega}$ of ${\llbracket}(\mathcal{M}_{1},\{(q,B)\}{\rrbracket}$ such that $C_{0}={{\textsl{inf}}}_{\mathcal{M}_{1}}(u(v)^{\omega})$ and there is a pair $(q_{0},B_{0})$ in $\alpha_{2}$ such that $q_{0}\in f(C_{0})$ and $B_{0}\cap f(C_{0})=\emptyset$. We add the pair $(q_{0},Q\setminus f(C_{0}))$ to $\alpha^{\prime}$ and note that by the argument in the preceding paragraph, ${\llbracket}(\mathcal{M},\alpha^{\prime}){\rrbracket}\subseteq{\llbracket}\mathcal{R}{\rrbracket}$.
If ${\llbracket}(\mathcal{M}_{1},\{(q,B)\}{\rrbracket}\subseteq{\llbracket}(\mathcal{M},\alpha^{\prime}){\rrbracket}$ then $\alpha^{\prime}$ is the desired acceptance condition. If not, there exists a word $u(v)^{\omega}$ such that for $C={{\textsl{inf}}}_{\mathcal{M}_{1}}(u(v)^{\omega})$ we have $q\in C$ and $C\cap B=\emptyset$, but either $q_{0}\not\in f(C)$ or $f(C)\cap(Q\setminus f(C_{0}))\neq\emptyset$. Because $C_{0}$ is the maximum SCC of $\mathcal{M}_{1}$ containing $q$ and containing no element of $B$, we have $C\subseteq C_{0}$, so $f(C)\subseteq f(C_{0})$ and therefore $q_{0}\not\in f(C)$. Let $C_{1}$ be the maximum SCC $C$ of $\mathcal{M}_{1}$ such that $C\subseteq C_{0}$, $q\in C$, and $q_{0}\not\in f(C)$. This is not empty, so there is a word $u^{\prime}(v^{\prime})^{\omega}$ such that $C_{1}={{\textsl{inf}}}_{\mathcal{M}_{1}}(u^{\prime}(v^{\prime})^{\omega})$, which is in ${\llbracket}\mathcal{R}{\rrbracket}$ because it satisfies $(q,B)$. Thus there exists a pair $(q_{1},B_{1})$ in $\alpha_{2}$ that is satisfied by $u^{\prime}(v^{\prime})^{\omega}$, and we add the pair $(q_{1},Q\setminus f(C_{1}))$ to the acceptance condition $\alpha^{\prime}$. As above, we have ${\llbracket}(\mathcal{M},\alpha^{\prime}){\rrbracket}\subseteq{\llbracket}\mathcal{R}{\rrbracket}$.
@@
A DRA $\mathcal{R}=(\mathcal{M}_{1},\alpha_{1})$ in singleton normal form with $|\alpha_{1}|=m$.
If ${\llbracket}\mathcal{R}{\rrbracket}\in\mathbb{IRA}$ then return an IRA accepting ${\llbracket}\mathcal{R}{\rrbracket}$, else return “no”.
@@
Assume the input is a DRA $\mathcal{R}=(\mathcal{M}_{1},\alpha_{1})$ in singleton normal form with $|\alpha_{1}|=m$. The algorithm $\mathbf{TestInIR}$ computes the right congruence automaton $\mathcal{M}$ of ${\llbracket}\mathcal{R}{\rrbracket}$, which has $n$ states, at most the number of states of $\mathcal{R}$. The main loop of the algorithm is executed at most $mn^{3}$ times, and each execution makes calls to the inclusion and equivalence algorithms for DRAs, and runs in time polynomial in the size of $\mathcal{R}$, so the overall running time of $\mathbf{TestInIR}$ is polynomial in the size of $\mathcal{R}$.